Problem 1 Use the following balance of payments data for the Buro Area from the IMF to answer the questions for 2013 1. What is the Euro Areas total for capital and finance account balance? 2. What is the Euro Areas total for current, capital, and finance account? 3. Does the Euro Areas BOP balance? Euro Area Balance of Payments Assumptions(billions of US$) A. Current account balance B. Capital account balance C. Financial account balance D. Net errors and omissions E. Reserves and related ite ms 2010 201120122013 2216 17143054 7.3 27.9 12.2 -103.2 -213.7-3452 7.3 14 33.7 1.3 1.8 10.1 38.4 01 34.8
Answer #1

1. Euro Area's total for capital and finance accounts 2013

Capital account balance = 27.9 Billion USD

Financial account = -345.2 Billion USD

Total = 27.9 -345.2 = -317.3 Billion USD

2. Euro Area's total for current, capital and finance accounts 2013

Current account balance = 305.4 Billion

Total for current, capital and finance accounts = 305.4 +27.9-345.2 = -11.3 Billion USD

3. As we see the current, capital and finance accounts do bot balance out each other to zero (as shown in #2) . Hence the Euro Area's BOP does not balance.